OpenStudy (cutiekawdfi):

help please (: find the real root by graphing 29. x^2+3=x^3-5

OpenStudy (jdoe0001):

well \(\bf x^2+3=x^3-5\implies x^2+3-x^3+5=0 \\ \quad \\ -x^3+x^2+8=0\implies -x^3+x^2+8=y\impliedby \textit{graph that}\)

OpenStudy (jdoe0001):

notice, 0 = y that is, what are the values of "x", when y is 0, or what are the x-intercepts or in quadratic lingo, what are the zeros and thus the solutions

OpenStudy (jdoe0001):

are you mean to use a graphing calculator? if so.. plug that in :) if you need one, many out there available

OpenStudy (jdoe0001):

just to avoid confusion =) are you meant to use a graphing calculator? is what I meant =)